OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 1.04x + 3.14 = 7.98

OpenStudy (anonymous):

1.04x + 3.14 = 7.98 subtract 3.14 on both sides 1.04x + 3.14 - 3.14 = 7.98 - 3.14 1.04x = 4.84 divide both sides by 1.04 1.04x/1.04 = 4.84/1.04 x = 4.65